OpenStudy (anonymous):

B=3, C=8, Sin A = 1/4 What's the Area of ABC?

OpenStudy (anonymous):

we know that sin x = opposite/hyp so in this case sin A = 1/4 where the opposite = y = 1 and the hyp. = c = 4 now we have a right triangle so we use this formula: \[A = (bh)/2\] where h = y = 1 and b = x = (15)^1/2 so : \[A = (\sqrt{15})(1)/2 = 1.94 \] correct me if I'm wrong, I'm not sure of the answer though